4. Triangular Arbitrage. Assume the following quotes: Citibank quotes U.S dollars per pound at $1.5400/£, National Westminster quotes euro per pound at €1.6000/£, Deutschebank quotes dollars per euro at $0.9700/€. Calculate how a market trader at Citibank with $1,000,000 can make inter-market arbitrage profit.
Answer #1

Start with 1000000 dollars

Convert dollars to pounds from Citibank=1000000/1.54=649350.64935 pounds

Convert pounds to euros from National Westminster=649350.64935*1.6=1038961.03896 euros

Convert euros to dollars from Deutsche Bank=1038961.03896*0.97=1007792.20779 dollars

Profit=1007792.20779-1000000=7792.20779 dollars